### v1

# Flow-Aware Ellipsoidal Filtration for Persistent Homology of Recurrent Signals

Abstract
One common use of persistent homology is to explore the shape of point clouds, where points are assumed to be sampled from a geometric object. We propose a novel filtration, called ellipsoidal filtration, which assumes that point clouds are sampled from a dynamic smooth flow. Instead of creating topologies from point clouds at increasing scales using isotropic balls (e.g. Vietoris–Rips filtration), ellipsoidal filtration creates ellipsoids around points based on local flow variances, approximating the flow’s manifold as the scale increases. We show that constructing ellipsoidal neighbourhoods improves the denoising of recurrent signals and the estimation of recurrence times, especially when the data contain bottlenecks. Choosing ellipsoids according to the maximum persistence of the $H_{1}$ class provides a data-driven threshold for both denoising and recurrence-time estimation.
keywords:
Persistent Homology , Adaptive Filters , Recurrent Signals , Ellipsoidal Filtration
[…]
[a]organization=School of Computer Science, University of Birmingham,city=Birmingham, postcode=B15 2TT, country=UK,, email= obe851@student.bham.ac.uk
[…]
[b]organization=Electrics and Electronics Department, Turkish-German University,city=Istanbul, postcode=34820, country=Turkey,, email= cihan.katar@tau.edu.tr
[…]
[c]organization=School of Computer Science, University of Birmingham,city=Birmingham, postcode=B15 2TT, country=UK,, email= m.little.1@bham.ac.uk
1 Introduction
Recurrent signals exhibit repeated patterns that manifest as close returns in their state-space trajectories, a phenomenon common in natural systems Little et al. (2007); Vieten and Weich (2020). The geometry of these trajectories encodes key information about the underlying dynamics—particularly the structure of the attractor, which indicates whether the signal is periodic, quasi-periodic, or chaotic (Kantz and Schreiber, 2003; Strogatz, 2018). In practice, the state space is reconstructed from scalar time series via time-delay (sliding-window) embeddings (Takens, 1981) or obtained directly from multivariate sensor recordings. This state space yields a discrete, high-dimensional point cloud—a set of points equipped with a distance metric.
Persistent homology, a method from applied and computational topology, characterises the topology of such point clouds across multiple spatial scales (Edelsbrunner et al., 2002; Edelsbrunner and Harer, 2008). As the scale parameter increases, data are mapped to topological spaces called simplicial complexes. The first step, a filtration, defines how points are connected at each scale. Common choices include the alpha, Čech, and Vietoris–Rips filtrations. Among these, the Vietoris–Rips filtration is often preferred for its computational and memory efficiency: points are connected when balls centred at them intersect, typically using the Euclidean distance.
Time-series analysis with persistent homology largely relies on sliding-window embeddings to convert one-dimensional signals into high-dimensional point clouds. We focus on point-cloud applications in the literature; for further details readers can refer to (Ravishanker and Chen, 2019; Perea, 2019; Chazal and Michel, 2021). Perea and Harer (2015) established a theoretical link between the most persistent $H_{1}$ class, signal periodicity, and the choice of window length and embedding dimension. Gakhar and Perea (2023) further showed that sliding-window embeddings of quasi-periodic signals are dense in an $n$-torus, which can be characterised via persistent homology. Vejdemo-Johansson et al. (2015) used persistent cohomology to construct circular coordinates from sliding-window embeddings, enabling periodicity detection, dimensionality reduction, and denoising; this was later generalised to toroidal coordinates for quasi-periodic signals (Scoccola et al., 2023).
Although these methods address periodicity, they use Vietoris–Rips-based filtrations that consider only spatial relationships and treat points as samples from a static geometric object. Moreover, circular coordinates parameterise topological $S^{1}$ classes rather than directly counting cycles in time. By contrast, we adopt a close-returns perspective (Kantz and Schreiber, 2003; Little et al., 2007): replacing $n$-dimensional balls with flow-aware ellipsoids allows us to detect first returns and estimate recurrence times for each point, even when many cycles occur with rapidly varying frequency (see Section 5.4).
Fernández et al. (2023) proposed the Fermat distance as a density-aware intrinsic metric for constructing topological spaces from point clouds, which can be robust to noise. However, because it does not incorporate temporal information, we observe that it may fail to recover the main loop of the attractor in narrow bottlenecks adjacent to dense–sparse regions (Section 5.1).
Kališnik et al. (2024) presented work related in spirit to ours by introducing ellipsoidal neighbourhoods for persistent homology on static point clouds sampled from manifolds. Ellipsoids are oriented via local Principal Component Analysis (PCA) on $k$-nearest neighbours with a globally fixed aspect ratio, and simplices are added when ellipsoids intersect. Their setting is spatial rather than time-series; temporal flow and spatio–temporal structure are not modelled. Their focus is homology recovery and classification against Vietoris–Rips; by contrast, we address time-series data with a flow-aware design for denoising and recurrence-time estimation, using adaptive ellipsoids that follow the local flow variance.
Our method was developed in the search for an adaptive neighbourhood for topological denoising of quasi-periodic signals (Robinson, 2016), where prior work used $k$-NN. We explore neighbourhood choice in terms of both cardinality and geometry. The geometric neighbourhood needs to be adaptive to the local flow in recurrent dynamics. We therefore define spatio–temporal neighbourhoods and adapt ellipsoids to the local flow, evaluating denoising and recurrence-time estimation on synthetic and real accelerometer data.
In our earlier work (Eryilmaz et al., 2025), we introduced an adaptive ellipsoidal filtration that incorporates local flow information by shaping anisotropic neighbourhoods according to signal gradients, enabling flow-aware persistent homology for recurrent dynamics. However, the local neighbourhood used only temporal windowing, which was insufficient to capture the variance in close returns.
In the present study, we extend this framework in several ways. First, we refine the adaptive ellipsoid construction by defining spatio–temporal neighbourhoods that jointly consider local spatial proximity and short-term temporal evolution, allowing more accurate estimation of local flow geometry. Second, we apply the method to real accelerometer recordings of recurrent motion, validating its practical relevance. Third, we compare ellipsoidal filtrations with Vietoris–Rips and Fermat filtrations to evaluate differences in topological representation. Finally, we assess the utility of the resulting neighbourhoods for recurrence-time estimation, demonstrating improved performance over Vietoris–Rips-based approaches. The main contributions are:
  • We show that ellipsoidal filtration outperforms Vietoris–Rips and Fermat filtrations in capturing the main loop of the attractor under bottlenecks and varying sampling density (Section 5.1).
  • We compare neighbourhood selection strategies for topological low-pass filtering (Robinson, 2016), evaluating neighbourhood size and geometry across signal-to-noise ratios. Ellipsoidal, flow-aligned neighbourhoods reduce attenuation of low-amplitude components (Sections 5.2–5.3).
  • We revisit recurrence period density methods (Little et al., 2007): replacing isotropic balls with ellipsoids derived from the adaptive filtration, and using persistent homology to justify scale selection, improves first-return and recurrence-time estimation (Section 5.4).
The paper is structured as follows. Section 2 reviews background on recurrent signals and persistent homology. Section 3 formulates the problem for temporally sampled dynamical data with observational noise and motivates flow-aware neighbourhoods. Section 4 presents the ellipsoidal filtration, our proposed method. Section 5 reports experiments on dense–sparse phase-space structure, denoising of synthetic and 3D accelerometer signals, and recurrence-time estimation. Section 6 discusses implications and limitations, and Section 7 concludes with future directions.
2 Background
2.1 Recurrent Signals
Recurrent signals are characterised by repeated patterns, which appear as returns to previously visited regions in state space. These patterns often reflect the presence of underlying deterministic dynamics.
We model the evolution of such signals as a continuous flow governed by a dynamical system. Let $\mathbf{u}(t)\in\mathbb{R}^{d}$ denote the latent state at time $t$, evolving according to
$\frac{d\mathbf{u}}{dt}=f(\mathbf{u}(t),t),$ (1)
where $f:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$ is a smooth, nonlinear function. Given an initial condition, the solution traces a trajectory in state space that represents the system’s temporal evolution.
In practice, we observe only a finite sequence of noisy, discrete-time samples:
$\mathbf{v}_{i}=\mathbf{u}(t_{i})+\mathbf{n}(t_{i}),\quad\text{with }t_{i}=t_{0}+i\Delta t,$ (2)
where $\mathbf{n}(t_{i})$ denotes additive noise. The resulting point cloud approximates the underlying trajectory and enables geometric and topological analysis.
We say that the system exhibits a recurrence at time $t$ if there exists a delay $\delta_{t}>0$ such that
$\mathbf{u}(t+\delta_{t})\in\mathbf{B}(\mathbf{u}(t),r),\quad\text{and}\quad\mathbf{u}(t+\tau)\notin\mathbf{B}(\mathbf{u}(t),r)\quad\text{for all }0<\tau<\delta_{t},$ (3)
where $\mathbf{B}(\mathbf{u}(t),r)$ is the closed ball of radius $r>0$ centred at $\mathbf{u}(t)$. The value $\delta_{t}$ is interpreted as the approximate recurrence time, indicating the system’s first return to a neighbourhood of its earlier state (Kantz and Schreiber, 2003; Little et al., 2007).
Recurrence time $\delta_{t}$ depends strongly on the neighbourhood radius $r$. If $r$ is too large, nearly all states may appear recurrent; if too small, most true recurrences may be missed. Choosing a suitable threshold is therefore essential, and various strategies have been proposed in the literature (Schinkel et al., 2008; Li et al., 2025).
This formulation naturally includes both periodic and aperiodic dynamics despite these two seems mutually exclusive. In the case of exact periodicity, recurrence times $\delta_{t}$ remain constant, and recurrence occurs at fixed intervals even as $r\to 0$. In contrast, real-world signals are generally aperiodic, and $\delta_{t}$ varies over time. In such cases, recurrence is approximate and highly dependent on the choice of neighbourhood size. The study of these variations—known as recurrence time statistics—has a long history in nonlinear time-series analysis (Kantz and Schreiber, 2003).
Understanding and analysing these time-varying recurrence structures requires methods that are robust to noise and sensitive to local geometry. Persistent homology provides a principled, multi-scale framework for this purpose, motivating our use of topological methods in the study of recurrent signals. In this study, we propose a data-driven approach that defines ellipsoidal neighbourhoods to determine the recurrence threshold based on local variance. The scale of these neighbourhoods is determined using persistent homology to capture the global topological structure of the signal.
2.2 Persistent Homology
Persistent homology is a foundational technique in topological data analysis (TDA) (Carlsson, 2009; Edelsbrunner and Harer, 2008). It identifies and tracks topological features—connected components, loops, and higher-dimensional voids—across a scale parameter $\varepsilon$.
The workflow has two steps. First, construct a filtration, i.e. a nested family of complexes $\{K_{\varepsilon}\}_{\varepsilon\geq 0}$ built from the data by a proximity rule (e.g. Vietoris–Rips: add a simplex when all pairwise distances of its vertices are $\leq\varepsilon$). As $\varepsilon$ increases, the complex becomes more connected and features appear and disappear. Second, compute homology on each $K_{\varepsilon}$ and record for every feature its birth scale $b$ and death scale $d$. The persistence (or lifetime) is $\ell=d-b,$ which quantifies the feature’s prominence across scales.
Formally, for a finite sample $\mathbb{X}\subset\mathbb{R}^{n}$, a filtration
$K_{\varepsilon_{1}}\subseteq K_{\varepsilon_{2}}\subseteq\cdots\subseteq K_{\varepsilon_{m}}$ (4)
is obtained by increasing $\varepsilon$. Exact recovery of the topology of the underlying space is often impossible with finite, noisy data, but persistent homology yields a robust multi-scale summary.
Figure 1: Illustration with a Vietoris–Rips (VR) filtration: (a) data and connections at increasing $\varepsilon$; (b) barcode showing $H_{0}$ (components) and $H_{1}$ (loops) as horizontal intervals; (c) persistence diagram (PD), mapping each feature to its birth–death point $(b,d)$. The dominant $H_{1}$ class is farthest from the diagonal.
3 Problem Statement
Point clouds are often modelled as samples from a distribution concentrated near a geometric object $X\subset\mathbb{R}^{n}$, with noise controlling the spread around $X$ (Carlsson, 2009). Under this view, common filtration techniques such as Vietoris–Rips and Čech are used to infer the Betti numbers of $X$ from a finite sample $\mathbb{X}$ (see Section 2.2). Because these constructions relate points solely through pairwise spatial distances, they impose a uniform, isotropic threshold to define neighbourhoods and build simplicial complexes. However, in point clouds derived from time-series data, the relationships between samples reflect both spatial and temporal proximity.
In this work, we consider a different setting in which the observed point cloud $\mathbb{V}\subset\mathbb{R}^{n}$ is sampled at uniform time intervals from a dynamical process with observation noise. We assume the existence of an unobserved latent trajectory $\mathbf{u}(t)$ evolving according to the same smooth flow introduced in Section 2.1, governed by Eq. (1), and we observe noisy samples given by Eq. (2).
Thus, the point cloud $\mathbb{V}=\{\mathbf{v}(t_{i})\}_{i=1}^{N}$ represents a time-parametrised dynamical process rather than a static geometric object, with its spatial density reflecting the local speed of the flow—denser in slow regions and sparser in fast ones. This anisotropic sampling invalidates the assumption that uniform spatial distances between points define meaningful neighbourhoods. Consequently, standard isotropic filtrations may fail to recover the true topological structure of the trajectory $\mathbf{u}(t)$.
In this paper, we address the following core problems, all centred on neighbourhood construction that respects the local dynamics of the flow:
  1. 1.
    designing a flow-aware filtration that approximates the manifold of the underlying trajectory using locally estimated, PCA-aligned ellipsoids instead of isotropic balls;
  2. 2.
    determining a data-driven scale, based on the death time of the dominant $H_{1}$ class, that captures the global structure while preserving local details;
  3. 3.
    performing topology-preserving denoising by averaging points within adaptive ellipsoidal neighbourhoods without distorting waveform morphology;
  4. 4.
    improving recurrence-time estimation by identifying first-return times through these adaptive, topologically consistent neighbourhoods.
4 Methods
This section describes how the proposed ellipsoidal filtration is constructed to incorporate local flow information from recurrent signals. The goal is to replace isotropic neighbourhoods, which ignore the direction and speed of evolution, with anisotropic ellipsoids that adapt to the local geometry of the trajectory. The construction proceeds in four main steps: (i) estimating local covariance from spatio–temporal neighbourhoods, (ii) defining adaptive ellipsoids aligned with local flow directions, (iii) detecting intersections between ellipsoids to determine connectivity, and (iv) constructing a simplicial complex suitable for persistent homology.
4.1 Local Covariance Estimation
The first step is to estimate how the signal varies locally around each sample. Given a point cloud representation of the signal $X=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}\subset\mathbb{R}^{d}$, each point $\mathbf{x}_{i}$ represents the system’s state at a discrete time step. To capture both the temporal evolution and the recurrence structure, we construct two types of local neighbourhoods:
  1. 1.
    an index-based (temporal) neighbourhood, which captures short-term evolution in time, and
  2. 2.
    a state-space (spatial) neighbourhood, which identifies recurrent states that are close in phase space but possibly far apart in time.
The temporal neighbourhood is defined by a symmetric window around each point:
$\mathcal{T}_{i}=\{\mathbf{x}_{j}\mid|j-i|\leq\tau\},$ (5)
where $\tau\in\mathbb{N}$ is the half-width of the window. This neighbourhood describes the short segment of the trajectory immediately surrounding the point $\mathbf{x}_{i}$.
The spatial neighbourhood consists of the $k$-nearest neighbours in Euclidean distance:
$\mathcal{S}_{i}=\text{$k$-nearest neighbours of }\mathbf{x}_{i}\text{ in state space}.$ (6)
This captures recurrent states that reappear near the same region of the attractor at possibly very different times.
The combined neighbourhood is given by
$\mathcal{N}_{i}=\mathcal{T}_{i}\cup\mathcal{S}_{i},$ (7)
which unites local temporal continuity with spatial recurrence. This union is crucial because it balances short-term temporal information with the broader geometric structure of the attractor.
The local geometry around each point is characterised by the empirical covariance matrix
$\Sigma_{i}=\frac{1}{|\mathcal{N}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{N}_{i}}(\mathbf{x}_{j}-\mathbf{x}_{i})(\mathbf{x}_{j}-\mathbf{x}_{i})^{\top}.$ (8)
This covariance matrix encodes both the direction and magnitude of local variation, effectively capturing how the trajectory is oriented and stretched in phase space. The eigenvectors of $\Sigma_{i}$ indicate the dominant flow directions, while the eigenvalues reflect how rapidly the signal varies along each direction.
4.2 Adaptive Ellipsoids from Local Covariance
Once the local covariance is obtained, it is used to construct an anisotropic neighbourhood in the form of an ellipsoid. Let
$\Sigma_{i}=Q_{i}\Lambda_{i}Q_{i}^{\top}$ (9)
be the eigendecomposition of the covariance matrix, where $Q_{i}$ contains the orthonormal eigenvectors and $\Lambda_{i}=\mathrm{diag}(\lambda_{i,1},\dots,\lambda_{i,d})$ the eigenvalues, ordered such that $\lambda_{i,1}\geq\cdots\geq\lambda_{i,d}>0$.
The adaptive ellipsoid centred at $\mathbf{x}_{i}$ and scale $\varepsilon>0$ is defined by
$E_{i}(\varepsilon)=\left\{\mathbf{x}\in\mathbb{R}^{d}\;\middle|\;(\mathbf{x}-\mathbf{x}_{i})^{\top}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{x}_{i})\leq\varepsilon^{2}\right\}.$ (10)
Each ellipsoid is aligned with the local flow directions (eigenvectors) and stretched according to the local variance (eigenvalues). The semi-axis lengths are proportional to $\varepsilon\sqrt{\lambda_{i,j}}$, meaning that as $\varepsilon$ increases, each ellipsoid expands uniformly while preserving its anisotropy. This allows the filtration to respect the local geometry of the trajectory at every scale.
4.3 Intersection of Adaptive Ellipsoids
To construct a topological complex, we must determine whether two ellipsoids intersect. If two ellipsoids overlap at scale $\varepsilon$, the corresponding points are connected by an edge.
Let $A=\Sigma_{i}/\varepsilon^{2}$, $B=\Sigma_{j}/\varepsilon^{2}$, and $v=\mathbf{x}_{i}-\mathbf{x}_{j}$. Following the formulation in Gilitschenski and Hanebeck (2012), define
$K(S)=1-v^{\top}\left(\frac{1}{1-S}B^{-1}+\frac{1}{S}A^{-1}\right)^{-1}v,\quad S\in(0,1).$ (11)
The two ellipsoids $E_{i}(\varepsilon)$ and $E_{j}(\varepsilon)$ intersect if and only if
$\min_{S\in(0,1)}K(S)\geq 0.$ (12)
Since $K(S)$ is convex on $(0,1)$, this minimisation is a one-dimensional convex optimisation problem. In practice, we employ a golden-section search with an early termination criterion: if a negative value of $K(S)$ is encountered during the search, the two ellipsoids are declared non-intersecting and the algorithm immediately proceeds to the next pair.
4.4 Ellipsoid Complex
Given the family of ellipsoids $\{E_{i}(\varepsilon)\}_{i=1}^{n}$ centred at the points of the dataset, we define the ellipsoid complex at scale $\varepsilon$ as
$\mathcal{K}_{\varepsilon}=\left\{\sigma\subseteq X\;\middle|\;E_{i}(\varepsilon)\cap E_{j}(\varepsilon)\neq\emptyset\ \text{for all distinct}\ i,j\in\sigma\right\}.$ (13)
Intuitively, two points are connected if their local uncertainty regions—represented by ellipsoids—overlap. Higher-order simplices are added whenever all pairwise intersections among their vertices are non-empty. This definition yields a flag complex whose 1-skeleton (graph) is determined by ellipsoid intersections.
As the scale parameter $\varepsilon$ increases, the ellipsoids grow and the complex becomes more connected, generating a nested sequence $\mathcal{K}_{\varepsilon_{1}}\subseteq\mathcal{K}_{\varepsilon_{2}}\subseteq\cdots,$ which defines a valid filtration for persistent homology analysis. In this way, ellipsoidal filtration generalises the Vietoris–Rips construction by replacing isotropic balls with anisotropic, flow-aware ellipsoids.
4.5 Algorithm for Building the Ellipsoid Complex
Algorithm 1 summarises the computational procedure for building the ellipsoid complex at a given scale. For each pair of points, the algorithm checks whether their adaptive ellipsoids intersect according to the convex criterion in Eq. (11). The resulting edge list defines the 1-skeleton, from which the flag complex is constructed. The flag complex construction and persistent homology were implemented with the SimplexTree data structure from the GUDHI library (Maria et al., 2014).
Algorithm 1 Ellipsoid Complex Construction at Scale $\varepsilon$
1:Point cloud $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$, covariance matrices $\{\Sigma_{i}\}$, scale $\varepsilon$
2:Initialise edge set $E\leftarrow\emptyset$
3:for $i=1$ to $n$ do
4:  for $j=i+1$ to $n$ do
5:   $A\leftarrow\Sigma_{i}/\varepsilon^{2}$,  $B\leftarrow\Sigma_{j}/\varepsilon^{2}$,  $v\leftarrow\mathbf{x}_{j}-\mathbf{x}_{i}$
6:   Compute $\min_{S\in(0,1)}K(S)$ using golden-section search with early termination
7:   if $\min K(S)\geq 0$ then
8:     Add edge $(i,j)$ to $E$
9:   end if
10:  end for
11:end for
12:Form the flag complex $\mathcal{K}_{\varepsilon}$ from the 1-skeleton $(X,E)$
13:return $\mathcal{K}_{\varepsilon}$
Repeating this process over an increasing sequence of scales $0<\varepsilon_{1}<\varepsilon_{2}<\cdots$ produces the filtration $\{\mathcal{K}_{\varepsilon_{t}}\}$ used for persistent homology computation. This filtration captures how topological features of the data—such as loops and connected components—emerge and disappear as neighbourhoods expand along the flow’s geometry.
Figure 2: Ellipsoids fitted to local neighbourhoods. Each neighbourhood includes the temporal window of three past and three future points, together with the 15 nearest spatial neighbours. The ellipsoids align with the local geometry estimated from these combined spatio–temporal neighbourhoods.
Code availability. The implementation and analysis scripts are (to be) hosted at github.com/omerberyilmaz/flow-aware-ellipsoidal-filtration.
5 Experiments and Results
5.1 Filtration behaviour with dense and sparse phase–space regions
We generate a two-dimensional trajectory from a Hamiltonian system with a non-linear potential, discretised using the Störmer–Verlet method:
$\dot{q}=p,\qquad\dot{p}=-\big(q^{r}-q-\epsilon\omega\sin(\omega q)\big),$ (14)
with parameters $T=17.22$, $h=0.08$, $r=5$, $\epsilon=0.39$, $\omega=3$, and initial conditions $(q_{0},p_{0})=(0.5,1.0)$. Gaussian noise is added independently to both coordinates to mimic measurement uncertainty. The resulting trajectory is shown in phase space $(p,q)$ in Figure 3.
Although the trajectory is sampled at uniform time intervals, the induced point cloud in phase space is highly non-uniform: slow dynamics yield dense regions, whereas fast transitions leave sparse ones. Because temporal information is ignored, a purely spatial approach cannot distinguish whether nearby points belong to the same dynamical phase. This limitation makes the system a useful test case for evaluating how different filtrations recover the trajectory’s topology from its point cloud representation. Importantly, the same fast–slow structure also arises in practice, for example in 3D accelerometer recordings of cyclic human motion.
We compare three filtrations: ellipsoidal, Vietoris–Rips (VR), and Fermat distance. VR depends only on Euclidean distance, whereas the Fermat distance incorporates sampling density through an intrinsic path-based metric (Fernández et al., 2023). For a finite set of points $\mathbb{X}_{n}$, the sample Fermat distance between $x,y\in\mathbb{X}_{n}$ is defined as
$d_{\mathbb{X}_{n},p}(x,y)=\inf_{\gamma}\sum_{i=0}^{r}\lvert x_{i+1}-x_{i}\rvert^{p},$ (15)
where the infimum ranges over all discrete paths $\gamma=(x_{0},x_{1},\dots,x_{r+1})$ with endpoints $x_{0}=x$ and $x_{r+1}=y$, and intermediate points $x_{1},\dots,x_{r}\in\mathbb{X}_{n}$. We set $p=2$ in this experiment.
Figure 3 compares the three filtrations across increasing neighbourhood scales. The aim is to recover the global loop that reflects the system’s recurrence, despite the narrow bottleneck in the middle. With Fermat distance (Figure 3a), dense regions connect quickly, but the bottleneck is bridged before sparse regions are included, preventing loop recovery. Vietoris–Rips (Figure 3b) also connects across the bottleneck too early and likewise fails to recover the loop. In contrast, the ellipsoidal filtration (Figure 3c) adapts neighbourhoods to local flow direction and speed. This allows the loop to appear at smaller scales while the bottleneck remains separated in the leftmost plot, giving a more accurate representation of the underlying topology.
These results show that neighbourhoods defined only by density or distance cannot recover meaningful topology, because the scale parameter does not reflect the local flow structure. Ellipsoidal filtration combines temporal information with spatial proximity, enabling the correct loop structure to be recovered even in the presence of bottlenecks. Although this dataset is synthetic, similar structures are common in real-world signals, such as accelerometer data from running and cycling (see Section 5.3). The following experiments build on this result, showing how appropriate neighbourhood selection is crucial for denoising recurrent signals and estimating recurrence times.
(a) Fermat
(b) Vietoris–Rips
(c) Ellipsoidal
Figure 3: Comparison of three filtrations on the same dataset: (a) Fermat distance, (b) Vietoris–Rips, and (c) Ellipsoidal filtration. Each panel shows a $1\times 4$ grid across increasing scale values.
5.2 Denoising a recurrent signal
We evaluate whether flow-aware neighbourhoods improve the denoising of recurrent signals. Building on earlier evaluations of topological filters (Eryilmaz et al., 2025), we replicate the experimental setting and make two changes: (i) a refined ellipsoidal filtration based on spatio–temporal PCA, and (ii) an updated synthetic dataset with the chirp phase defined as the integral of the instantaneous frequency, $\phi(t)=2\pi\!\int_{0}^{t}f(s)\,ds$.
The synthetic signal is designed to expose two representative challenges: a linearly increasing frequency and a narrow bottleneck in the trajectory. We define
$x(t)=10\cos\!\big(\phi(t)\big),\qquad y(t)=2\sin\!\big(\phi(t)\big)\,S\!\big(x(t)\big),\qquad t\in[0,2],$ (16)
with $n=500$ uniform samples. The instantaneous frequency is
$f(t)=f_{\mathrm{start}}+\frac{f_{\mathrm{end}}-f_{\mathrm{start}}}{t_{\max}}\,t,\quad(f_{\mathrm{start}},f_{\mathrm{end}})=(1,10)\ \text{Hz},\quad t_{\max}=2,$ (17)
and the chirp phase is
$\phi(t)=2\pi\!\int_{0}^{t}f(s)\,ds=2\pi\!\left(f_{\mathrm{start}}t+\frac{(f_{\mathrm{end}}-f_{\mathrm{start}})}{2t_{\max}}t^{2}\right).$ (18)
Anisotropy in $y$ is introduced by
$S(x)=1-0.9\exp\!\left(-\tfrac{1}{2}\left(\frac{x}{\max_{t}|x(t)|}\right)^{2}\right).$ (19)
We add independent, zero-mean Gaussian noise to each coordinate, with axis-wise variances $\sigma_{x}^{2}=P_{x}/\mathrm{SNR}$ and $\sigma_{y}^{2}=P_{y}/\mathrm{SNR}$, where $P_{x}$ and $P_{y}$ are the mean-square powers of the clean $x$ and $y$ signals. This yields a bivariate Gaussian noise with diagonal covariance, horizontally elongated $(\sigma_{x}>\sigma_{y})$ and temporally white; the overall 2D SNR equals the target value. Under severe noise, the most persistent $H_{1}$ scale can be unreliable; here we therefore treat the $H_{1}$ death time as a guide rather than a guarantee.
We compare five strategies: (i) a Vietoris–Rips-based (isotropic, “spherical”) topological filter, (ii) an ellipsoidal (flow-aware) topological filter built using the spatio–temporal PCA construction as described in Section 4, (iii) a $k$-NN topological filter with $k=20$, (iv) a fixed 20-sample moving average, and (v) an adaptive moving average whose window length is set from an FFT-based frequency estimate (Nyquist rule). For Vietoris–Rips and ellipsoidal topological filters, the neighbourhood scale is chosen by the death time of the most persistent $H_{1}$ class.
Even when the noise level is modest (20 dB), the methods differ markedly in how well they preserve the input structure (Figure 4). Panel (a) shows the noisy input with a clear frequency rise and a narrow low-amplitude bottleneck.
In panel (b), the fixed 20-sample moving average is the most clearly destructive: once the period drops below the window length, the filter averages across opposite lobes and the waveform is driven towards zero in the middle segment; fine peak details disappear. In panel (c), the adaptive moving average is not adaptive enough to preserve the signal structure. As the frequency increases, it fails to adjust its window size accordingly, leading to reduced peak-to-peak amplitude, rounded peaks, and loss of fine details.
Topological filters behave differently for the low-amplitude component (see Fig. 4(b)). The $k$-NN filter in panel (d) preserves oscillations over large portions of the record, yet in the bottleneck its fixed-size neighbourhood spans opposing directions of the trajectory, causing local attenuation and shape distortion. The VR (spherical) filter in panel (e) is more severely affected because isotropic balls cut across the thin geometry; cross-bottleneck averaging reduces peak-to-peak amplitude in the low-amplitude region. By contrast, the ellipsoidal filter in panel (f) aligns neighbourhoods with the local flow direction and speed, avoiding cross-bottleneck mixing; amplitude and phase are preserved across the record, including the bottleneck.
Figure 5 reports root-mean-square error (RMSE) versus SNR for $x(t)$ and $y(t)$. For $x(t)$, the moving-average baselines can yield lower RMSE at low SNRs because aggressive smoothing suppresses noise; however, their error quickly plateaus as SNR increases. The three topological filters improve steadily with SNR and overtake the baselines from moderate SNRs upwards, with broadly similar performance among the topological methods on $x(t)$. For $y(t)$, the contrast is stronger: the ellipsoidal filter achieves the lowest RMSE over the mid- to high-SNR range, while $k$-NN and spherical VR also surpass the time-domain baselines once the SNR is moderate. This behaviour reflects a trade-off: time-domain smoothing reduces noise but also attenuates peak-to-peak structure, so its RMSE gains saturate, whereas topology-based neighbourhoods preserve structure and therefore continue to improve as noise decreases.
(a) Filtered outputs for the high-amplitude component $x(t)$.
(b) Filtered outputs for the high-amplitude component $y(t)$.
Figure 4: Comparison of filters applied to denoise the signal $x(t)$ and $y(t)$ at SNR = 20 dB. (a) Original and noisy signals, (b) Moving average,(c) Adaptive moving average, (d) $k$-NN topological filter, (e) Spherical topological filter, (f) Ellipsoidal topological filter.
(a) RMSE vs. SNR for the high-amplitude component $x(t)$.
(b) RMSE vs. SNR for the low-amplitude component $y(t)$.
Figure 5: Quantitative comparison of filters across SNR levels. Ellipsoidal filtration achieves the lowest RMSE, particularly for the low-amplitude component $y(t)$.
5.3 3D accelerometer data denoising
We now evaluate the method on real data: publicly available tri-axial accelerometer recordings from cycling (Vieten and Weich, 2019). The data were sampled at 500 Hz; we analysed a 3,500-sample segment from the first participant.
To set comparable neighbourhood scales for Vietoris–Rips (VR) and the ellipsoidal filtration, we first computed persistent homology on a shorter 350-sample segment (about one full cycle). Let $B$ and $D$ be the birth and death of the dominant $H_{1}$ class, and L=DBL=D-B its lifetime. For each filtration we examined four scales: $B$, $B+\tfrac{1}{2}L$, $D$, and $D+\tfrac{1}{2}L$. When estimating ellipsoidal covariances, we pooled neighbours across adjacent cycles to stabilise local variance estimates, as described in the Methods.
Figure 6 shows persistence diagrams for the same 3D segment under the Vietoris-Rips and the proposed Ellipsoidal filtrations (red: $H_{0}$, blue: $H_{1}$). In the VR diagram, two $H_{1}$ classes are clearly more persistent, whereas in the ellipsoidal diagram a single $H_{1}$ class dominates. The blue points farthest from the diagonal represent the most persistent loop in each panel. In the denoising and recurrence time estimation 5.4 we consider the scale of the most persistent $H_{1}$ class.
Figure 7 compares spherical and ellipsoidal neighbourhood averaging across four scales. At small scales, both methods preserve the data’s loop structure. As the scale increases, spherical neighbourhoods span opposite sides of narrow passages and average across bottlenecks, compressing the geometry and attenuating amplitude. In contrast, ellipsoidal averaging follows the local flow and better preserves loops across scales.
Figure 8 applies denoising by taking the geometric median of the samples within each sphere or ellipsoid. The geometric median yields smoother reconstructions and exhibits greater robustness to noise than the mean.
(a) VR filtration persistence diagram
(b) Ellipsoidal filtration persistence diagram
Figure 6: Persistence diagrams computed on the 3D accelerometer signal. (a) Vietoris–Rips (spherical) filtration. (b) Ellipsoidal filtration.
(a) s=2.494,e=2.052s=2.494,\ e=2.052
(b) s=3.626,e=3.334s=3.626,\ e=3.334
(c) s=4.759,e=4.616s=4.759,\ e=4.616
(d) s=5.891,e=5.898s=5.891,\ e=5.898
Figure 7: Accelerometer denoising using mean averaging: comparison of spherical versus ellipsoidal neighbourhoods at increasing scales. For each filtration, the scale is set relative to the most persistent $H_{1}$ feature: $b$, $\tfrac{1}{2}\ell, b+b+\ell$, and $b+\tfrac{3}{2}\ell$, where $b$ is the birth scale and $\ell$ the lifetime. The corresponding numerical values are shown in subcaptions. The left column shows the raw data, the middle column the spherical neighbourhood outcomes, and the right column the ellipsoidal neighbourhood outcomes.
(a) s=2.494,e=1.967s=2.494,\ e=1.967
(b) s=3.626,e=3.419s=3.626,\ e=3.419
(c) s=4.759,e=4.872s=4.759,\ e=4.872
(d) s=5.891,e=6.325s=5.891,\ e=6.325
Figure 8: Accelerometer denoising using geometric median averaging: comparison of spherical versus ellipsoidal neighbourhoods at increasing scales. For each filtration, the scale is set relative to the most persistent $H_{1}$ feature: $b$, $\tfrac{1}{2}\ell, b+b+\ell$, and $b+\tfrac{3}{2}\ell$, where $b$ is the birth scale and $\ell$ the lifetime. The corresponding numerical values are shown in subcaptions. The left column shows the raw data, the middle column the spherical neighbourhood outcomes, and the right column the ellipsoidal neighbourhood outcomes.
Figure 9: Accelerometer time-series at scale corresponding to the most persistent $H_{1}$ feature (s=5.891,e=5.898s=5.891,\ e=5.898). The three panels show the $x$-, $y$-, and $z$-axis accelerations respectively.
5.4 Recurrence-time estimation
We assess whether the same flow-aware neighbourhoods that improved denoising (Sections 5.25.3) also improve recurrence-time estimation. The concept of recurrence and first-return times was introduced in Section 2.1. For consistency, we use the chirp signal defined in Section 5.2, which combines a linearly increasing frequency, anisotropic deformation, and a narrow bottleneck as described earlier.
In the experimental setting, we consider a discrete sequence of noisy samples $\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$. At each index $i$, we define a neighbourhood $\mathcal{N}_{i}\subset\mathbb{R}^{d}$ either as a closed Euclidean ball
$\mathcal{N}_{i}=\{\,x\in\mathbb{R}^{d}:\|x-x_{i}\|_{2}\leq\varepsilon\,\}$ (20)
(spherical case), or as a locally aligned ellipsoid
$\mathcal{N}_{i}=\{\,x\in\mathbb{R}^{d}:(x-x_{i})^{\top}A_{i}(x-x_{i})\leq 1\,\},$ (21)
where $A_{i}\in\mathbb{R}^{d\times d}$ is a positive-definite matrix estimated from local PCA (ellipsoidal case).
The first recurrence time at index $i$ is then defined as
$T_{1}(i)=\min\{\,j>i:j-i\geq\tau_{\min},\ x_{j}\in\mathcal{N}_{i},\ \text{and}\ x_{k}\notin\mathcal{N}_{i}\ \ \forall\,i<k<j\,\},$ (22)
where $\tau_{\min}=15$ samples prevents spurious early returns introduced by noise or oversampling. For reference, ground-truth first returns are computed from the unwrapped phase $\theta(t)$ as the smallest $j>i$ such that $\theta[j]\geq\theta[i]+2\pi$. As in Section 5.3, we evaluate four scales per filtration taken from the dominant $H_{1}$ class: $B$, $B+\tfrac{1}{2}(D-B)$, $D$, and $D+\tfrac{1}{2}(D-B)$.
Figures 11 and 12 show the recurrence structure across scales. On the clean signal, ellipsoidal neighbourhoods detect almost all first returns across scales except at the birth scale. In contrast, fixed-radius balls struggle as the period shortens: first returns are frequently missed at all scales, and at larger scales balls span opposite sides of the bottleneck, producing spurious early returns. Under noise, ellipsoidal neighbourhoods capture more first returns, missing only one first return at the largest scale. Importantly, no spurious early returns are caused by the bottleneck which makes it possible to estimate returns for all points.
(a) Clean — Vietoris–Rips
(b) Clean — Ellipsoidal
(c) Noisy — Vietoris–Rips
(d) Noisy — Ellipsoidal
Figure 10: Persistent diagrams for clean and noisy signals under two filtrations. Columns compare Vietoris–Rips (left) to ellipsoidal (right); rows compare clean (top) to noisy (bottom).
(a) $\varepsilon=0.223$
(b) $\varepsilon=0.519$
(c) $\varepsilon=0.815$
(d) $\varepsilon=1.112$
(e) $r=0.240$
(f) $r=0.305$
(g) $r=0.371$
(h) $r=0.437$
Figure 11: Clean signal: recurrence structure under ellipsoidal filtration (top, increasing $\varepsilon$ left to right) and Vietoris–Rips filtration (bottom, increasing $r$ left to right).
(a) $\varepsilon=0.223$
(b) $\varepsilon=0.445$
(c) $\varepsilon=0.667$
(d) $\varepsilon=1.112$
(e) $r=0.210$
(f) $r=0.284$
(g) $r=0.359$
(h) $r=0.434$
Figure 12: Noisy signal: recurrence structure under ellipsoidal filtration (top, increasing $\varepsilon$ left to right) and Vietoris–Rips filtration (bottom, increasing $r$ left to right).
6 Discussion
Across all experiments, adaptive ellipsoidal neighbourhoods captured the shape of the trajectory, even in data with changing frequencies and a narrow bottleneck. In the Hamiltonian system, they recovered the shape of attractor at smaller scales without joining the bottleneck too early, whereas Vietoris–Rips and Fermat filtrations connected dense regions prematurely and failed to recover the attractor shape (Fig. 3). In the synthetic chirp, they preserved both amplitude and signal shape through the bottleneck and gave the most accurate reconstruction, particularly for the small-amplitude part $y(t)$ (Figs. 4 and 5). On the 3D accelerometer data, they revealed one clear loop that remained stable across scales, while spherical averaging caused the structure to collapse at larger scales (Figs. 6 and 7).
These outcomes follow from how neighbourhoods are constructed. As described in Section 4, we assume that the point clouds are sampled from a flow. We then determine the local flow region as the union of spatial $k$-nearest neighbours and a short temporal window. PCA is applied to this union of points to estimate the variance in the local flow. By aligning ellipsoids with the principal components and stretching them according to the variances, we approximate the flow using ellipsoids as the scale increases. In contrast, the Vietoris–Rips and Fermat filtrations cannot incorporate this flow model.
In denoising (Section 5.2), all topological filters—where we examined the size of the neighbourhoods in terms of both cardinality ($k$-NN) and geometry (spherical and ellipsoidal) improved with SNR, but adaptive ellipsoids best preserved amplitude and phase in the bottleneck, yielding lower RMSE, especially for $y(t)$ (Figs. 4 and 5).
On 3D accelerometer signals (Section 5.3), ellipsoidal filtration maintained the loop structure across the four scales associated with the dominant $H_{1}$ class, whereas spherical averaging increasingly compressed the narrow passages (Figs. 6 and 7). Using the geometric median instead of the mean made the method more robust to outliers and revealed a clearer attractor structure (Fig. 8).
For recurrence-time estimation (Section 5.4), scales derived from the most persistent $H_{1}$ class tie the recurrence threshold to the intrinsic geometry of the trajectory rather than to an arbitrary radius. Recurrence times were then obtained from the first-return intervals, as described in Section 2.1. Using these scales, the flow-aligned neighbourhoods detected first returns robustly under strong frequency drift and avoided spurious early returns across the bottleneck (Figs. 11 and 12).
Since the dominant $H_{1}$ feature can shorten towards the diagonal or be replaced by spurious loops at low $\mathrm{SNR}$, selecting the denoising or recurrence threshold solely from the most persistent $H_{1}$ class may be unreliable. In our experiments (medium–high $\mathrm{SNR}$) with finite sample lengths, extreme excursions were uncommon and the dominant $H_{1}$ provided a stable guide; nevertheless, at lower $\mathrm{SNR}$ we caution against relying on a single $H_{1}$-based scale without additional checks.
Compared with the Vietoris–Rips filtration, the ellipsoidal approach introduces two additional parameters that control the number of spatial and temporal neighbours. The union of these neighbours defines the local region used to estimate the local flow geometry via PCA. The resulting eigenvectors indicate the dominant flow directions, while the eigenvalues determine how the ellipsoid is stretched along each axis. Although the ellipsoid is later scaled beyond this region during the filtration, its orientation and shape are derived from the local dynamics. This additional parametrisation improves adaptability but increases computational cost, since constructing the spatial tree and performing PCA for each point are computationally intensive steps.
In a broad sense, VR and ellipsoidal filtrations both consider entropy, in the context of uncertainty, as the scale parameter used to connect points. While VR creates isotropic (spherical) uncertainty regions around each point and increases these with the scale parameter, ellipsoidal filtration creates anisotropic (ellipsoidal) uncertainty regions. Notably, ellipsoidal filtration produces smaller ellipsoids in slow-flow regions and larger ones in fast-flow regions. In this way, it evaluates how the topology evolves over time without explicitly using time as a parameter but instead through entropy. When the uncertainty regions of two points intersect, we assume that, at that scale, they are no longer separable and therefore connect them to form edges and higher-dimensional simplices. An analogy can be drawn between ellipsoidal filtration and motion blur photography, as illustrated in Fig. 13
(a) Motion blur with temporal structure.
(b) Blur without temporal awareness.
Figure 13: Analogy between ellipsoidal filtration and motion blur photography. (a) Motion blur with temporal structure (Image credit: Matti Sulanto, used with permission). (b) Blur without temporal awareness (Image credit: İnsunur Eryılmaz, used with permission).
7 Conclusion and Future Work
We propose a flow-aware ellipsoidal filtration that treats the observed point cloud as samples from a smooth flow. Using spatio–temporal neighbourhoods and local covariance (PCA), it fits ellipsoids aligned with the local direction and speed, yielding an adaptive anisotropic filtration. Future work includes comprehensive tests on real data sets and applications in higher-dimensional settings.